**Please answer ASAP

You are responsible for managing a portfolio of two investments for the coming year. You have read all of the latest financial news and according to your favourite reputable financial analysts, the probability of another recession occurring is 30% and the probability of a boom is 20%; otherwise the economy will be behaving normally. Common shares of Firm A are expected to earn a return of 30% during boom times, but lose 10% during a recession. Normally, the firm’s common shares would earn a 20% return. The price of Firm B’s common shares is expected to decline 10% during boom times, but increase by 6% during a recession. Normally the firm’s common shares would earn a 9% return.

Questions:

G) Compare the return versus risk of a portfolio 100% invested in Firm B's shares with the return versus risk of the portfolio consisting of 30% invested in Firm A and 70% in Firm B. Which would you prefer? Why?                                                                                     

h. Find the Minimum Variance Portfolio (MVP) consisting of some combination of shares of both Firm A and Firm B. What proportion of the portfolio must be invested in Firm A's common shares and what proportion in those of firm B to achieve the MVP? (Use the "Solver" tool in Excel.)                                                                                     

I) Calculate the standard deviation of expected returns on the MVP?                                                                                     

J) Calculate the expected return on the MVP?                                                                                  
Answer #1
COMMON SHARES OF FIRM A
A B C=A*B D=B-13 E=D^2 F=E*A
Scenario Probability Return Return *probability Deviation Deviation Deviation squared*Probability
Percent fromexpected Squared
Recession 0.3 -10 -3 -23 529 158.7
Normal 0.5 20 10 7 49 24.5
Boom 0.2 30 6 17 289 57.8
SUM 13 SUM 241
Expected Return 13 percent
Variance 241
Standard Deviation 15.52 Percent (Square Root of 241)
COMMON SHARES OF FIRM B
A B C=A*B D=B-4.3 E=D^2 F=E*A
Scenario Probability Return Return *probability Deviation Deviation Deviation squared*Probability
Percent fromexpected Squared
Recession 0.3 6 1.8 1.7 2.89 0.867
Normal 0.5 9 4.5 4.7 22.09 11.045
Boom 0.2 -10 -2 -14.3 204.49 40.898
SUM 4.3 SUM 52.81
Expected Return 4.3 percent
Variance 52.81
Standard Deviation 7.27 Percent (Square Root of 52.81)
Return Risk
Share of Firm A 13 15.52
Share of Firm B 4.3 7.27
Correlation of A&B
Scenario Return A Return B
Recession -10 6
Normal 20 9
Boom 30 -10
CORRELATION -0.58004 (USING "CORREL "function of Excel)

Firm A Shares Assetl Firm B Shares-Asset2 Return of asset 1-R1-13% Return of asset):R2-4.3% Standard deviation of asset 1-s1-15.52% Standarc deviation of asset 2:S2-7 27% Correlation of asset 1 and 2-Corr(1,2)0.58 Covariance(1,2)- Corr(,2)S1S2 0.58* 15.527.2765.4416 W1-Investment in asset 1 (FirmA)-0.3 w2-Investment inasset 2(Firm B):0.7 Porttolio Return 3 13+0.7*4.3 6.91% 6.91 Equation (1) Variance-(w12)(S1 2)+(w2 2)(S22)+2w1w2t Cov(1,2) Portfolio Variance(w12)15.522)+w22)7.27 2)+2w1w2 (-654416) Portfolio Variance-(w1A2)24 1 +(W2A2752 81-w1w2*130.88 w1 lnvestment in FirmA 0.3, W2 0.7 Equation (2) Portfolio Variance (0.32)2410.7 2)*52,81-0.30,7* 130.88 Portfolio Variance Deviation-Square root of 20.0821 Portfolio Standard Deviation Square root (20.0821) 4.48% 4.481305613 Risk(Sta ndard Deviatio Return Portfolio withOLDTIO Variance:- Deviation-Square root OT 20.0821 Portfolio Standard Deviation Square root (20.0821) 4.48% 4481305613 RiskfSta ndard Deviatio Return n) Portfolio with 100% FirmB Portfoliowith 30% in A and 70% in B 4.30%) 7.27% 6.91% 4.48% ALL POSSIBLE PORTFOLIOS Vp(Using Equation (2) Rp w1 13+w2 4.3 Weight Weight of of FirmA |Firm B Portfolio |Return(96) Portfolio Variance 52.81 0.9 5.17Weight of of FirmA Firm B Return%) Portfolio Portfolio Variance 52.81 0.9 0.85 0.8 0.75 0.7 0.65 0.6 0.55 0.5 0.45 0.4 0.35 0.3 0.25 0.2 0.16 5.17 0.15 0.2 0.25 0.3 0.35 0.4 0.45 0.5 0.55 0.6 0.65 5.605 26.890525 6.04 6.475 20.228125 20.0821 7.345 22.059525 8.21532.384725 40.7325 9.085 51.203725 63.7984 9.955 78.516525 6.91 7.78 8.65 9.52 10.39 0.75 0.8 0.84 0.85 0.87 0.9 0.95 10.825 114.323125 11.26 135.4116 11.608 153.811264 0.13 11.869 168.502861 12.13 183.9589 12.565 211.417725 0.050.2 11.26 135.4116 1.608153.811264 11.695 158.623525 11.869 168.502861 12.13 183.9589 12.565211.417725 241 0.84 0.85 0.15 0.13 0.95 0.05 LOWEST VARIANCE PORTFOLIO nvestment in Firm A nvestment in Firm B Variance 30% 70% Portfolio Expected Return Portfolio Standard Deviation